OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through (0,6) and is parallel to the line described by y=2x+3.

OpenStudy (luigi0210):

Take the slope from the equation of \(m=2\) and the points \((0,6)\) and plug it into \(y-y_{1}=m(x-x_{1})\)

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Well you got \(m\) and \((x_{1},~y_{1})\), just plug them in and solve.

OpenStudy (luigi0210):

The number on \(y_{1}\) is just a subscript not an exponent. And you got it backwards :/

OpenStudy (anonymous):

ok..

OpenStudy (luigi0210):

We have the point-slope formula: \(\large y-y_{1}=m(x-x_{1})\) And we have the coordinates: \(\large (0, 6)==>(x_{1},~y_{1})\) And the slope: \(\large m=2\) from \(\large y=2x+3~===>~y=mx+b\) where \(m=slope\)

OpenStudy (luigi0210):

So plugging that in.. \(\large y-(6)=2(x-0)\) Can you simplify that into \(\large y=mx+b\)?

OpenStudy (anonymous):

Yes. Thanks for the help.

OpenStudy (luigi0210):

No problem, and good luck!